OpenStudy (anonymous):

Solve by elimination. 2x − y = 5 3x − 2y = 6 (0, negative-3) (3, 4) (4, 3) (1, 4) @iGreen

OpenStudy (igreen):

2x - y = 5 3x - 2y = 6 Take the first one and rearrange it: 2x - y = 5 Subtract 2x to both sides: -y = -2x + 5 Multiply -1 to all terms: y = 2x - 5 Plug in '2x - 5' into the second equation for y: 3x - 2y = 6 3x - 2(2x - 5) = 6 Distribute 2 into the parenthesis: 3x - 4x + 10 = 6 Subtract: -x + 10 = 6 Subtract 10 to both sides: -x = -4 Multiply -1 to all terms: x = 4 Now that we have the x-value, we can plug that into any of the two equations to find the y-value. 2x - y = 5 Plug in x = 4: 2(4) - y = 5 Multiply: 8 - y = 5 Subtract 8 to both sides: -y = -3 Multiply -1 to all terms: y = 3 So the answer is (4, 3)

OpenStudy (anonymous):

Solve by elimination. 8x + y = -16 -3x + y = -5 (-1, -8) No solution (8, 1) (-8, 1)